# Impact of Large Language Model Assistance on Patients Reading Clinical Notes: A Mixed-Methods Study

### v1

@@ Appendix E Supplemental File 3: Interview Questions
Next, the researcher shared a custom link with the participant for them to access the user interface and view the clinical note they previously submitted. The researcher prompts the participant to refamiliarize themselves with the original note and to share any thoughts they have out loud. After about five minutes, the researcher asks the following questions
@@
The participant is then directed to the options menu, where they can enable and read descriptions of each augmentation. The participant is prompted to read about each augmentation and enable the one that most interests them. The participant is then asked to explore their clinical note with the active augmentation and to share their thoughts out loud. The researcher then asks the following questions
@@
Then the participant is prompted to return to the Options menu, disable the active augmentation, and repeat the process with the next augmentation they are interested in.
Once all of the augmentations have been explored individually, the participant is prompted to enable as many augmentations as they would like, to explain the selection they made, and provide overall feedback on their experience reading their unmodified and modified notes.
Appendix F Supplemental Table: Common themes from qualitative interviews
Current practices and attitudes towards note reading
Reading notes to gather information
“I wanted to know the dates when that was diagnosed, and maybe when was the last time I did a test” [P7]
“Figure out when my appointments are. I’ll look at blood work, you know, if anything’s weird.” [P5]
Reading notes to check the narrative
“[To find out] what my doc actually thought as opposed to what he said to me” [P4]
“You kind of want that insight into like what the doctor’s take was on it” [P2]
Medical terminology makes notes difficult to understand
“Usually, I have to Google some words. Especially in the beginning when you’re first diagnosed, there’s a lot of initials and words that a lot of people don’t understand” [P1]
“At the time, I looked them up but they’re hard to retain even. It’s just I don’t use them in everyday language” [P2]
“I have to do a lot of Googling though to see what this means and what that means and whether it’s significant or not.” [P3]
Wins with the modifications
Improved understanding of the note
“It gives me more comfort and information going, oh, that’s what they mean” [P6]
“It gives you all the information and it’s actually less stressful to read. I think because it’s not as overwhelming.” [P2]
“It’s definitely easier to understand and at least one of the questions I had, I got answered here” [P5]
Ease of use
“This is like a personalized Wikipedia page for my diagnosis” [P3]
“It’s just nice to have it all there if I want it. I don’t have to look up everything but I can look up what I want to.” [P2]
Concerns with the modifications
Missing or misleading information
“I feel like this makes assumptions that can’t really be made” [P5]
“I guess I feel like blood work should be in there.” [P6]
More information can be confusing
“It made me feel a little bit more worried because they were things that I hadn’t conceived of yet” [P2]
“That’s a bit overly heavy on medical jargon.” [P3]
“If I’m going to have a medication list, I think I would want the medication list of what I’m on or what I will be having, not everything that came up.” [P5]
Appendix G Supplemental Figure 5: Full codebook for coding interviews
@@
Accuracy; check if they got facts right
“I want to have an understanding of what’s going on and see if anything has been overlooked.” [P3]
Alignment
Check if doctors got the context/story right
Check if doctor/patient priorities align
“To make sure that the doctor and I had the same conversation. Did I miss anything?” [P1]
Memory; to remember something or see if they missed anything that was discussed
“…if I wanted to review my notes, you know, post a procedure, this is very helpful to recall what was done and what it all meant.” [P7]
Preparation; to review and ready questions for the visit
“Sometimes the notes will reflect the test I just had. If I get them before our meeting I like to look at them so I can ask them questions.” [P1]
Comprehension; to check your own understanding
“I like to understand …the definitions of what I’m reading.” [P7]
Information-seeking
I consider myself detail-oriented and like more information
More information can be comforting
“Sometimes research is how I self-soothe. The more information I have, it makes me feel more prepared and equipped to deal with any scenario.” [P2]
Space; to digest the visit alone and at their own pace
“During the meeting I understood everything but when you’re in the meeting you’re taking it all in and you really don’t know which questions to ask at the time and you kind of have to think afterwards.” [P1]
To understand the doctor’s perspective or thinking
“You kind of want that insight into …what the doctor’s take was on it.” [P2]
For health monitoring
Get results
Check appointments
“Wanting to see …trends and being able to really toggle back and forth between past …results and past things.” [P2]
@@
Medical terminology can make the note hard to understand
“It’s like having to translate [the note] from a different language.” [P2]“I have to do a lot of Googling though to see what this means and what that means and whether it’s significant or not.” [P3]
Fear due to a lack of understanding
“As I was reading it, I was wondering how bad it was. There was the term invasive. And that was scary.” [P3]
Patient portal usability
“My medical place uses MyChart and sometimes it’s not the friendliest to go back and forth in-between…” [P2]
Difficulty finding answers through self research
“I would look for the answer, which was extremely time consuming, and I didn’t always get a good answer.” [P2]“You have no idea how many papers I have…I wouldn’t have to flip through everything just to look for [one] definition or one appointment.” [P4]
@@
Translating medical terminology helps with understanding
“If I want to do more research I can but now I have a little snapshot of what they’re talking about.” [P1]
Convenient to see relevant information and answers in one place
“I think it’s helpful. Instead of having to run to Google.” [P1]
Simplified notes can seem more personal
“Like a patient took these notes, not a machine.” [P7]
This modification makes the note easier to read
Less stressful
“[The simplified note] is actually less stressful to read. I think because it’s not as overwhelming.” [P2]
Prompt new ideas and questions
“It also asked questions that I wouldn’t think to ask, that I would like the information for.” [P4]
Keeps patients focused
“I wouldn’t go off on all of these different websites…this doesn’t give me the worst possible scenario, I like that.” [P4]
Patients can do learning on their own time
“I could go home and look at them whenever, or, you know, however long it took me to understand it…” [P4]
Mod can help caregivers
“Cancer patients are not are not always–I couldn’t even read at that time, or pay bills–I was a hot mess. So I like the fact that the caretakers can take care of [the to do list].” [P4]
Modification helps with memory or organization
“I usually kind of look at the clinical notes just to make sure that I didn’t miss what they told me when we were in the exam.” [P6]
Different levels of simplification are needed at different times by different patients
“I like that it’s almost like you can increase how complicated it is as you feel ready for it.” [P2]“I would imagine there are definitely people who would want the more simplified version.” [P5]
Tool has intuitive performance
“I like how it phrased things, how it chose to simplify things.” [P2]“This presents some more information that I didn’t have and it’s detailed but it’s not too detailed.” [P3]
Being able to see the original version is important
Builds trust in the AI system
Allows the patient to compare what the doctor wrote
I would click on the simplified [note] just to get..bullet points. If I really wanted to have more detailed information, I would click on the original.” [P7]
@@
The modifications may miss something important
“”I feel like there’s a tiny bit of anxiety being like, ‘I’m having to trust that the AI got it right and that this is correct.’” [P2]“If you simplify it too much, people are going to miss some things and be like, ‘Why didn’t anyone tell me this? Why didn’t anyone tell me that?’.” [P5]
The modifications may lose context in their simplification
“I don’t know that it 100% captures the doctor’s thinking in the same way.” [P5]“I know what kind of treatment I had but if someone else were to read this, what is ‘extra cancer treatment’? Was it radiation? Was it chemo?” [P1]
An oversimplification may mislead patients
“I think it would make people think they understood more than they may actually understand.”
Professionalism and respect for gravity of situation
“I almost feel like it’s kind of talking down to you. It reads to me like at a third grade level.” [P6]“It’s so impersonal that this feels a little casual but it is a lot easier to understand.” [P2]
Too much information could be concerning/worrisome
“It almost adds more questions because it’s like, what is the difference between calcifications and punctate calcifications?” [P2]
Introduces more complexity (e.g., definitions with jargon)
“Some words I still don’t understand in the definitions.” [P4]“I don’t understand…prophylactic antibiotics. That’s–I don’t necessarily know what that means, so I would have to go back and look it up.” [P4]
The use of generic versus brand names can be confusing
“The inconsistency with the naming…[I could] have gotten the wrong information from this…” [P5]
Mod is not useful
“It doesn’t really add anything” [P7]“I’m not a fan of a to-do list here. Just because: check off ‘continue to monitor symptoms’. You don’t want to check that off.” [P1]
Errors the patient identified in the mods
“I assume ‘LMP’ means last menstrual period. I don’t know why it says over here that it was 12/17 and here it says it was 11/17. That doesn’t make sense.” [P5]“It’s a little weird…we went from this [AC-T] to this [T-AC] and it’s not clear that those two things are the same thing.” [P5]
@@
A sense of “betrayal”
“My body had turned traitor, so they could all be rebelling and I wouldn’t know. So it was that sense of surprise and not trusting my body.” [P2]
Anxiety related to prognosis
“Basically the big fear was, ‘Am I going to have chemotherapy and radiation?’ and ‘Is this terminal?’ I just sort of had blinders on to, ‘Where is this headed?’” [P2]
Patients learn more about their disease over time
“…the earlier you are in your treatment, the more helpful this will be. I’m several years of doing this…” [P1]
@@
“I would personally like the definitions to be closer. Instead of in the middle.” [P2]
Appendix H Supplemental Figure 6: Categories used for errors
Type of error
Example modification classified as an error
Reasoning behind error
Major errors
Hallucinated information
Common Question
@@ Appendix H Supplemental Figure 6: Categories used for errors
What are the pros and cons of having the mastectomy and tissue expander placement in Location0 versus Location1? Having the mastectomy and tissue expander placement in Location0 might be more convenient if you live there, but it requires follow up visits. On the other hand, doing the procedure in Location1 would allow you to return there for tissue reconstruction when needed, but may involve more logistical challenges.
Modification mixed up the description of Location0 versus Location1
Incorrect medical reasoning
Common Question
@@
How successful was my bilateral mastectomy and SLNB in addressing my invasive ductal carcinoma? The bilateral mastectomy and SLNB were performed to remove the cancerous tissue and check if it has spread to nearby lymph nodes. Your report indicates that no lymph nodes were involved (N0), which is a positive outcome as it suggests that the cancer might not have spread to other areas.
N0 is a description of the stage of the breast cancer, and not a reflection of the success of the surgery.
Incorrect medical fact or expansion
Simplified
@@
The patient had surgery on their right breast to remove a type of breast cancer called invasive ductal carcinoma with some features of lobular cancer.
Breast cancer can have lobular features. This augmentation implies that there was a separate type of cancer, lobular cancer.
Omission in augmentation
Simplified
The patient is a 47-year-old woman who has had treatment for a specific type of breast cancer (DCIS ER/PR negative HER2+).
@@
The original noted the staging of the breast cancer as pT1c, which is critical information to include even in a simplified version of a note
Phrasing could affect patient
Definition
vus POLE: A variant of uncertain significance (VUS) in the POLE gene, which may or may not contribute to cancer development
@@
Recommend changing to: “A variant of uncertain significance (VUS) in the POLE gene, which indicates we do not have enough information to understand whether this genetic mutation contributes to cancer development” in order to minimize worry
Time/tense issue
Structured
Repeat Echo on 11/19
@@
The original note stated “Echo 11/28=70%-2/7=60%-repeat 11/19=65%” which indicates the 11/19 echo was already completed
Minor errors
Assumed context
Definition
Lower outer quadrant: The lower outer section of the breast, divided into four parts called quadrants
@@
Lower outer quadrant can refer to a specific part of many different types of tissues, not just breast. This definition may lead patients to believe lower outer quadrant is a terminology specific to breast tissue.
Did not simplify
Simplified
@@
Get blood tests (CBCD, CMP) one year before returning
Recommend simplifying the abbreviations CBCD and CMP
Needs generic name
Definition
@@
Adjuvant TH: Adjuvant therapy given after surgery or radiation, combining taxol (a chemotherapy drug) and herceptin (targeted therapy for HER2-positive breast cancer)
Recommend providing the generic names for Taxol and Herceptin
Phrasing is awkward
Common Question
What is adjuvant hormonal therapy and why do you strongly recommend it for my treatment plan?
@@
The phrasing of “why do you” may make the reader think the clinician is specifically answering this question
Provider preference
Definition
Left breast ductal carcinoma in situ
Ductal carcinoma in situ
@@
The definition for “left breast ductal carcinoma in situ” is redundant if “ductal carcinoma in situ” has already been defined
Unnecessary excess information
Definition
@@
Nipple discharge: fluid seeping out of the nipple, which may be a symptom of a benign disorder or a sign of breast cancer
“which may be a symptom…” is unnecessary commentary on the definition
Appendix I Supplementary File 7: Web Survey Details
@@
How confident are you at filling out medical forms by yourself?
@@
Participants were given the following instruction ”Before we begin the study, we want to understand how much you already know about cancer. Please answer these questions WITHOUT looking at any other resources or websites.”.

### v2

@@ Appendix B Interview Questions
Next, the researcher shared a custom link with the participant for them to access the user interface and view the clinical note they previously submitted. The researcher prompted the participant to refamiliarize themselves with the original note and to share any thoughts they have out loud. After about five minutes, the researcher asked the following questions
@@
The participant was then directed to the options menu, where they could enable and read descriptions of each augmentation. The participant was prompted to read about each augmentation and enable the one that most interested them. The participant was then asked to explore their clinical note with the active augmentation and to share their thoughts out loud. The researcher then asked the following questions
@@
Then the participant was prompted to return to the Options menu, disable the active augmentation, and repeat the process with the next augmentation they were interested in.
Once all of the augmentations were explored individually, the participant was prompted to enable as many augmentations as they would like, to explain the selection they made, and provide overall feedback on their experience reading their original and augmented notes.
Appendix C Full Codebook for Coding Interviews
@@
Accuracy; check if they got facts right “I want to have an understanding of what’s going on and see if anything has been overlooked.” [P3]
Alignment Check if doctors got the context/story right Check if doctor/patient priorities align “To make sure that the doctor and I had the same conversation. Did I miss anything?” [P1]
Memory; to remember something or see if they missed anything that was discussed “…if I wanted to review my notes, you know, post a procedure, this is very helpful to recall what was done and what it all meant.” [P7]
Preparation; to review and ready questions for the visit “Sometimes the notes will reflect the test I just had. If I get them before our meeting I like to look at them so I can ask them questions.” [P1]
Comprehension; to check your own understanding “I like to understand …the definitions of what I’m reading.” [P7]
Information-seeking I consider myself detail-oriented and like more information More information can be comforting “Sometimes research is how I self-soothe. The more information I have, it makes me feel more prepared and equipped to deal with any scenario.” [P2]
Space; to digest the visit alone and at their own pace “During the meeting I understood everything but when you’re in the meeting you’re taking it all in and you really don’t know which questions to ask at the time and you kind of have to think afterwards.” [P1]
To understand the doctor’s perspective or thinking “You kind of want that insight into …what the doctor’s take was on it.” [P2]
For health monitoring Get results Check appointments “Wanting to see …trends and being able to really toggle back and forth between past …results and past things.” [P2]
@@
Medical terminology can make the note hard to understand “It’s like having to translate [the note] from a different language.” [P2]“I have to do a lot of Googling though to see what this means and what that means and whether it’s significant or not.” [P3]
Fear due to a lack of understanding “As I was reading it, I was wondering how bad it was. There was the term invasive. And that was scary.” [P3]
Patient portal usability “My medical place uses MyChart and sometimes it’s not the friendliest to go back and forth in-between…” [P2]
Difficulty finding answers through self research “I would look for the answer, which was extremely time consuming, and I didn’t always get a good answer.” [P2]“You have no idea how many papers I have…I wouldn’t have to flip through everything just to look for [one] definition or one appointment.” [P4]
@@
Translating medical terminology helps with understanding “If I want to do more research I can but now I have a little snapshot of what they’re talking about.” [P1]
Convenient to see relevant information and answers in one place “I think it’s helpful. Instead of having to run to Google.” [P1]
Simplified notes can seem more personal “Like a patient took these notes, not a machine.” [P7]
This augmentation makes the note easier to read Less stressful “[The simplified note] is actually less stressful to read. I think because it’s not as overwhelming.” [P2]
Prompt new ideas and questions “It also asked questions that I wouldn’t think to ask, that I would like the information for.” [P4]
Keeps patients focused “I wouldn’t go off on all of these different websites…this doesn’t give me the worst possible scenario, I like that.” [P4]
Patients can do learning on their own time “I could go home and look at them whenever, or, you know, however long it took me to understand it…” [P4]
Mod can help caregivers “Cancer patients are not are not always–I couldn’t even read at that time, or pay bills–I was a hot mess. So I like the fact that the caretakers can take care of [the to do list].” [P4]
Augmentation helps with memory or organization “I usually kind of look at the clinical notes just to make sure that I didn’t miss what they told me when we were in the exam.” [P6]
Different levels of simplification are needed at different times by different patients “I like that it’s almost like you can increase how complicated it is as you feel ready for it.” [P2]“I would imagine there are definitely people who would want the more simplified version.” [P5]
Tool has intuitive performance “I like how it phrased things, how it chose to simplify things.” [P2]“This presents some more information that I didn’t have and it’s detailed but it’s not too detailed.” [P3]
Being able to see the original version is important Builds trust in the AI system Allows the patient to compare what the doctor wrote I would click on the simplified [note] just to get..bullet points. If I really wanted to have more detailed information, I would click on the original.” [P7]
@@
The augmentations may miss something important “”I feel like there’s a tiny bit of anxiety being like, ‘I’m having to trust that the AI got it right and that this is correct.’” [P2]“If you simplify it too much, people are going to miss some things and be like, ‘Why didn’t anyone tell me this? Why didn’t anyone tell me that?’.” [P5]
The augmentations may lose context in their simplification “I don’t know that it 100% captures the doctor’s thinking in the same way.” [P5]“I know what kind of treatment I had but if someone else were to read this, what is ‘extra cancer treatment’? Was it radiation? Was it chemo?” [P1]
An oversimplification may mislead patients “I think it would make people think they understood more than they may actually understand.”
Professionalism and respect for gravity of situation “I almost feel like it’s kind of talking down to you. It reads to me like at a third grade level.” [P6]“It’s so impersonal that this feels a little casual but it is a lot easier to understand.” [P2]
Too much information could be concerning/worrisome “It almost adds more questions because it’s like, what is the difference between calcifications and punctate calcifications?” [P2]
Introduces more complexity (e.g., definitions with jargon) “Some words I still don’t understand in the definitions.” [P4]“I don’t understand…prophylactic antibiotics. That’s–I don’t necessarily know what that means, so I would have to go back and look it up.” [P4]
The use of generic versus brand names can be confusing “The inconsistency with the naming…[I could] have gotten the wrong information from this…” [P5]
Mod is not useful “It doesn’t really add anything” [P7]“I’m not a fan of a to-do list here. Just because: check off ‘continue to monitor symptoms’. You don’t want to check that off.” [P1]
Errors the patient identified in the mods “I assume ‘LMP’ means last menstrual period. I don’t know why it says over here that it was 12/17 and here it says it was 11/17. That doesn’t make sense.” [P5]“It’s a little weird…we went from this [AC-T] to this [T-AC] and it’s not clear that those two things are the same thing.” [P5]
@@
A sense of “betrayal” “My body had turned traitor, so they could all be rebelling and I wouldn’t know. So it was that sense of surprise and not trusting my body.” [P2]
Anxiety related to prognosis “Basically the big fear was, ‘Am I going to have chemotherapy and radiation?’ and ‘Is this terminal?’ I just sort of had blinders on to, ‘Where is this headed?’” [P2]
Patients learn more about their disease over time “…the earlier you are in your treatment, the more helpful this will be. I’m several years of doing this…” [P1]
@@
Suggestions for the augmentations “I would personally like the definitions to be closer. Instead of in the middle.” [P2]
Appendix D Categories of Errors
Type of error Example augmentation classified as an error Reasoning behind error
@@ Appendix D Categories of Errors
Hallucinated information Common Question What are the pros and cons of having the mastectomy and tissue expander placement in Location0 versus Location1? Having the mastectomy and tissue expander placement in Location0 might be more convenient if you live there, but it requires follow up visits. On the other hand, doing the procedure in Location1 would allow you to return there for tissue reconstruction when needed, but may involve more logistical challenges. Augmentation mixed up the description of Location0 versus Location1
@@
Incorrect medical reasoning Common Question How successful was my bilateral mastectomy and SLNB in addressing my invasive ductal carcinoma? The bilateral mastectomy and SLNB were performed to remove the cancerous tissue and check if it has spread to nearby lymph nodes. Your report indicates that no lymph nodes were involved (N0), which is a positive outcome as it suggests that the cancer might not have spread to other areas. N0 is a description of the stage of the breast cancer, and not a reflection of the success of the surgery.
@@
Incorrect medical fact or expansion Simplified The patient had surgery on their right breast to remove a type of breast cancer called invasive ductal carcinoma with some features of lobular cancer. Breast cancer can have lobular features. This augmentation implies that there was a separate type of cancer, lobular cancer.
@@
Omission in augmentation Simplified The patient is a 47-year-old woman who has had treatment for a specific type of breast cancer (DCIS ER/PR negative HER2+). The original noted the staging of the breast cancer as pT1c, which is critical information to include even in a simplified version of a note
@@
Phrasing could affect patient Definition vus POLE: A variant of uncertain significance (VUS) in the POLE gene, which may or may not contribute to cancer development Recommend changing to: “A variant of uncertain significance (VUS) in the POLE gene, which indicates we do not have enough information to understand whether this genetic mutation contributes to cancer development” in order to minimize worry
@@
Time/tense issue Structured  Repeat Echo on 11/19 The original note stated “Echo 11/28=70%-2/7=60%-repeat 11/19=65%” which indicates the 11/19 echo was already completed
@@
Assumed context Definition Lower outer quadrant: The lower outer section of the breast, divided into four parts called quadrants The original note stated, “…lower outer quadrant of the left breast…” Lower outer quadrant can refer to a specific part of many different types of tissues, not just breast. This definition may lead patients to believe lower outer quadrant is a terminology specific to breast tissue.
@@
Did not simplify Simplified Get blood tests (CBCD, CMP) one year before returning Recommend simplifying the abbreviations CBCD and CMP
@@
Needs generic name Definition Adjuvant TH: Adjuvant therapy given after surgery or radiation, combining taxol (a chemotherapy drug) and herceptin (targeted therapy for HER2-positive breast cancer) Recommend providing the generic names for Taxol and Herceptin
@@
Phrasing is awkward Common Question What is adjuvant hormonal therapy and why do you strongly recommend it for my treatment plan? The phrasing of “why do you” may make the reader think the clinician is specifically answering this question
@@
Provider preference Definition Left breast ductal carcinoma in situ Ductal carcinoma in situ The definition for “left breast ductal carcinoma in situ” is redundant if “ductal carcinoma in situ” has already been defined
@@
Unnecessary excess information Definition Nipple discharge: fluid seeping out of the nipple, which may be a symptom of a benign disorder or a sign of breast cancer “which may be a symptom…” is unnecessary commentary on the definition
Appendix E Web Survey Details
@@
How confident are you at filling out medical forms by yourself? (Note for appendix: health literacy question [10])
@@
Participants were given the following instruction ”Before we begin the study, we want to understand how much you already know about cancer. Please answer these questions WITHOUT looking at any other resources or websites.”. These questions are from the CHLT-6 cancer literacy instrument [13].